OpenStudy (anonymous):

x minus 7 divide by 3 =10

OpenStudy (anonymous):

Your question is X-7÷3=10 ?

OpenStudy (anonymous):

are you asking x-7/3=10 or (x-7)/3=10?

OpenStudy (anonymous):

I think he meant the second but I am not sure...

OpenStudy (anonymous):

me niether :(

OpenStudy (anonymous):

haha yeah >.<